Use the Method of Distribution Functions 3. (5 marks) The joint density of X and Y is given as , else Use the Method of Distribution Functions Find the probability density function of U- X-Y.
Answer #1

The joint PDF is -X F(X, у) 3D е *; 0 <у <х .

The region of integration 0 U X- Y u = n>0 is shown shaded below.

Y х-у 3D0 х-у 3Dи X (и, 0)

The PDF of  U =X -Y| is found using the CDF method. We see that in the region X x > o, U BX-Y0

The CDF is

P(U u) P (X- Y<u) coo P (U u) e dydx + хр^рx ә X-u е * dydx Fu (u) + xрАрx ә X-u = Fu (u) Fu (u) [-xex[-e ueu Fu (u) 1 - eu;

The PDF is

α fu (u) Fu(u) du f (u) e;u>0

Thus, we see that U =X -Y| has exponential distribution.